OpenStudy (anonymous):

Please help!! I give medals!! The point Negative twelve comma sixteen. would not fall on the line for which of the following representations? Seven minus two y equals two x minus 1. two y minus twenty-eight equals negative four x plus twelve. Negative three y equals six x plus twenty-four. Ten y equals negative five x plus one hundred.

OpenStudy (anonymous):

(-12,16) 7-2y=2x-1 7-2(16)=2(-12)-1 7-32=-24-1 -25=-25 2y-28=-4x+12 2(16)-28=-4(-12)+12 32-28=48+12 4=60 -3y=6x+24 -3(16)=6(-12)+24 -48=-72+24 -48=-48 10y=5x+100 10(16)=-5(-12)+100 160=60+100 160=160

OpenStudy (anonymous):

thanks!!!